How can I use Minitab to solve b and c Steps please
12. Use Minitab to simulate taking a random sample of 200 farm-raised turkey NG0, 8) ). See below for instructions to do the simulation Do a histogram of your sample. Describe its shape. a. Histogram of Turkey 25 20 7.5 22.5 100 7.5 C1 The histogram is evenly spread What are the z-scores of the maximum and minimum values in your sample? (Remember to use population (that is, distribution) parameters (30 and 8) to standardize.) Do these surprise you? Why or why not? b. c. According to the empirical rule, how many of the values do you expect to be les than 14 or greater than 46? Briefly explain your answer. d. Use Minitab to standardize all of your random numbers, then do a stem-and-leaf
Answer #1

Following to code to solve part b and c

Part b

MTB > let x1=minimum(x)............where x is smaple of 200 from N(30,8)

MTB > let xn=max(x)

MTB > let z1= (x1-mean(x))/8.697       

MTB > let zn=(xn-mean(x))/8.697

Z1 and Z2 are z score of normal distribution and it is not surprising to get these values since it lies between range d standard normal from -3 to 3

Part c

MTB > let c6=sum(x<14)

MTB > let c7=sum(x>46)

from c6 and c7 we get total number values less than 14 and greater than 46